OpenStudy (anonymous):

Find the function value? f(x)=x^3 f(1)=

OpenStudy (anonymous):

That is 1*3 right? which would be 3

OpenStudy (anonymous):

f(1)=1^3 = 1*1*1=1

OpenStudy (anonymous):

The caret (^) means exponent. In other words, your function is raising 1 to the third power, which equals one.

OpenStudy (anonymous):

ok that make sence

OpenStudy (anonymous):

answer is 1